# The evolution of inharmonicity and noisiness in contemporary popular music

### v1

@@ Abstract
Much of Western classical music uses instruments based on acoustic resonance. Such instruments produce harmonic or quasi-harmonic sounds. On the other hand, since the early 1970s, popular music has largely been produced in the recording studio. As a result, popular music is not bound to be based on harmonic or quasi-harmonic sounds. In this study, we use modified MPEG-7 features to explore and characterise the way in which the use of noise and inharmonicity has evolved in popular music since 1961. We set this evolution in the context of other broad categories of music, including Western classical piano music, Western classical orchestral music, and musique concrète. We propose new features that allow us to distinguish between inharmonicity resulting from noise and inharmonicity resulting from interactions between relatively discrete partials. When the history of contemporary popular music is viewed through the lens of these new features, we find that the period since 1961 can be divided into three phases. From 1961 to 1972, there was a steady increase in inharmonicity but no significant increase in noise. From 1972 to 1986, both inharmonicity and noise increased. Then, since 1986, there has been a steady decrease in both inharmonicity and noise to today’s popular music which is significantly less noisy but more inharmonic than the music of the sixties. We relate these observed trends to the development of music production practice over the period and illustrate them with focused analyses of certain key artists and tracks.
@@ 1 Introduction
Most musical instruments that perform Western classical music are based on the acoustic resonance principle. Such instruments produce harmonic or quasi-harmonic complex tones. In contemporary popular music, the means of production are more diversified. Drums play a key role, and they do not typically produce quasi-harmonic complex tones. Instruments may be based on electronic or digital sound production techniques. The instruments’ outputs may be heavily processed with effects. Musicians may not even use instruments and produce the entirety of the music on a computer. As a result, given a contemporary popular music track, it is possible that the description of the corresponding signal using quasi-harmonic complex tones is unsuitable.
In the spectral domain, a quasi-harmonic complex tone may be described as a series of partials whose frequency values are near multiples of one frequency value (the fundamental). There are at least two ways audio content may diverge from being quasi-harmonic: as exemplified in Lavengood, (2017, p. 22), the energy may not be centered around partials, in which case the signal may be described as ‘noisy’; or, as exemplified in Lavengood, (2017, p. 24), the partial frequency values may not be harmonically related, in which case the tone may be described as ‘inharmonic’.
@@
One key conclusion of this study is that contemporary popular music is more inharmonic than the music from the three other datasets considered. Our diachronic analysis of the evolution of noisiness and inharmonicity in CPM leads to the conclusion that, between 1961 and 1986, recorded music evolved in a general direction that goes from acoustic, tonal instruments playing content from a musical score (or, at least, that can be transcribed to a score), to heavily processed, noisy, and inharmonic content made in the recording studio and including drums. In more recent decades, music has become slightly more harmonic.
The remainder of this paper is structured as follows. The core content of the study can be found in sections 3 and 4 (the ‘core sections’). Section 3 describes the datasets and the signal features used in the study. Section 4 presents the results of the analysis. Sections 5 to 11 elaborate on several notions introduced in the core sections. Section 5 reflects on the relationship between the means of production and the produced music. It elaborates on the reasoning that led to examining noise and inharmonicity in contemporary popular music. Section 6 details characteristics of the signal that influence the inharmonic character of the music, discusses the link between inharmonicity and acoustic beating (in other words the acoustical ‘grit’), and substantiates the definition of inharmonicity we use in the core sections. Section 7 details the reasoning leading to the definition of ‘noisiness’ we use in the core sections. Section 8 addresses the influence of the recording medium and remastering processes. Section 9 relates the results from the core sections to those of previous diachronic studies. Section 10 details the influence of the psychoacoustic weighting used in the core sections. Section 11 presents detailed analyses of specific music tracks from the perspective proposed in the core sections. The paper concludes by summarising the findings from all sections.
@@ 2 Background and motivation
A defining property of a harmonic tone is that it has a fundamental frequency or $f_{0}$. A classic music information retrieval task consists in estimating the fundamental frequency of a sound, often taken as a synonym for ‘pitch tracking’ or ‘pitch estimation’ (Kim et al.,, 2018; Riou et al.,, 2023). The models performing such tasks are trained on harmonic or quasi-harmonic audio. For instance, Riou et al., (2023) uses MIR-1K, a dataset of quasi-harmonic samples (Hsu and Jang,, 2009) and MDB-stem-synth, a dataset of ‘exact multiples of the $f_{0}$(Salamon et al.,, 2017).
@@
Another classic music information retrieval task is audio-to-MIDI alignment (Raffel and Ellis,, 2016), in which MIDI files are aligned to a corresponding audio content. MIDI files have been referred to as symbolic ‘versions’ of pieces of music (Ewert et al.,, 2012; Raffel and Ellis,, 2016), a ‘transcription’ (Raffel,, 2016; Turetsky and Ellis,, 2003; Benetos et al.,, 2018) of the pitched content of the music. Such content has been characterised as ‘harmonic’ (Ono et al.,, 2008) and likely to being played with ‘harmonic instruments’ (Ewert et al.,, 2014).
@@
In the pitch tracking and audio-to-midi alignment tasks, pitched content is generally considered harmonic or quasi-harmonic. Yet, interaction with actual professional music production may show a different picture. For example, Figure 1 shows the power spectrum of one bass ‘note’ in the 2023 track ‘¡Fire!’ by Primaal. [^1] https://www.youtube.com/watch?v=_rd3IeLsfh8 Primaal is a brand of the Hyper Music production company mentioned by Deruty et al., (2022). In 2022 and 2023, Primaal authored music for commercials commissioned by brands such as L’Oréal, [^2] https://www.youtube.com/watch?v=UdQFK_k3VvY Adidas, [^3] https://www.youtube.com/watch?v=WKWUJ--Y05A Vichy, [^4] https://www.youtube.com/watch?v=R-box0e2J_E Honda, [^5] https://www.youtube.com/watch?v=9YXR1c4fW1Q GoPro, [^6] https://www.youtube.com/watch?v=UZ5G0tVwBkI&t=341s and Chanel. [^7] https://www.youtube.com/watch?v=VllN0yINA5A Therefore, it is possible to consider that Primaal’s production reflects an existing trend in the music industry. We write ‘note’ in scare-quotes, as its content is not harmonic even though it provides an impression of pitch. This example is not unique. In Primaal’s production, many other parts provide an impression of pitch while not featuring harmonic content.
@@
Figure 1: Power spectrum values for the first bass ‘note’ in Primaal’s ‘¡Fire!’. The vertical red and blue lines represent the fundamental and expected positions of the harmonics for the main two spectral peaks.
@@
There, therefore, seems to be a discrepancy between such examples of the use of inharmonicity in real-world music production and what appears to be the standard view within music information retrieval that pitched content corresponds to harmonic tones. This discrepancy provides motivation for exploring the extent to which pitched elements in CPM are harmonic.
3 Methods
@@ 3.1 Datasets
    The popular music dataset, or ‘BEA dataset’, contains 30435 tracks released between 1961 and 2022. There are at least 460 tracks for each year. This dataset is an extension of the dataset used by Deruty and Pachet, (2015). The choice of tracks for the datasets stems from the ‘Best Ever Albums’ website, [^9] https://www.besteveralbums.com/ a review aggregator that proposes the best-rated albums for each year of production. The ratings are dynamic: new reviews may modify an album’s ranking, making the dataset a specific snapshot at the time of writing. The dataset indiscriminately involves original and remastered versions. The influence of the remastering process is discussed in section 8.
@@
    The orchestra dataset contains 10800 orchestral and opera tracks from various sources, including Deutsche Grammophon’s ‘Classical Gold’, ‘The History of Classical Music’, Decca’s ‘55 Great Vocal Recitals’ and ‘Ultimate Boxset’ series. Tracks in this dataset range from the Baroque era to the early 20th century. Again, different interpretations of the same original score may be found in the dataset.
@@
    The musique concrète dataset consists of 1000 tracks from composers related to Pierre Schaeffer’s school of thought by either having directly collaborated with him or having produced music at INA/GRM in Paris. It includes music by Pierre Henry, Bernard Parmegiani, Denis Dufour, and François Bayle.
@@ 3.2.1 HR-inharmonicity
To measure inharmonicity, we use the Matlab R2022b implementation of the MPEG-7 feature HarmonicRatio. [^10] https://fr.mathworks.com/help/audio/ref/harmonicratio.html HarmonicRatio has been described as measuring ‘the proportion of harmonic components within the power spectrum’ (ISO,, 2001, p. 36), and ‘the ratio of harmonic power to total power’ (Moreau et al.,, 2006, p. 33). To our knowledge, HarmonicRatio is the only feature in the literature that measures the proportion of harmonic components in the general case. Features such as Inharmonicity in Peeters, (2004, p.17) measure ‘the […] divergence from a purely harmonic signal’ and can only be measured in relation to a single complex tone. HarmonicRatio derives from the normalised auto-correlation of the signal. The normalisation is performed so that the auto-correlation at zero lag equals one. The feature output is the maximum value of the normalised auto-correlation after the first zero-crossing. HarmonicRatio assumes values between 0 and 1, with harmonic complex tones resulting in 1. As we want to measure inharmonicity, we use a feature called HR-inharmonicity, defined as $1-HarmonicRatio$. Section  6 elaborates on the relationship between different aspects of the signal and HR-inharmonicity.
Patterson et al., (1996) and Yost, (1996, 1997) have argued that the ‘first peak of the auto-correlation function’ can be used to evaluate the ‘saliency or the strength of pitch of complex sounds’—in other words, the perceptual impression of pitch strength. The matter was further investigated by Shofner and Selas, (2002). The illustrations in Yost, (1996, p. 3330) and Shofner and Selas, (2002, p. 439) show without ambiguity that the ‘first peak of the auto-correlation function’ is the maximum value of the normalised auto-correlation after the first zero-crossing—that is, what the MPEG-7 standard designates as HarmonicRatio (ISO,, 2001; Moreau et al.,, 2006). One of our main concerns in this study is with quantifying the degree to which a signal deviates from being harmonic. However, as HarmonicRatio can be used to gauge the strength of pitch, the properties of the signal that we find to result in lower (or higher) HarmonicRatio values may also lead to a corresponding decrease (or increase) in the perceived strength of pitch.

### v2

@@ Abstract
Much of Western classical music relies on instruments based on acoustic resonance, which produce harmonic or quasi-harmonic sounds. In contrast, since the mid-twentieth century, popular music has increasingly been produced in recording studios, where it is not bound by the constraints of harmonic sounds. In this study, we use modified MPEG-7 features to explore and characterise the evolution of noise and inharmonicity in popular music since 1961. We place this evolution in the context of other broad categories of music, including Western classical piano music, orchestral music, and musique concrète. We introduce new features that distinguish between inharmonicity caused by noise and that resulting from interactions between discrete partials. Our analysis reveals that the history of popular music since 1961 can be divided into three phases. From 1961 to 1972, inharmonicity in popular music, initially only slightly higher than in orchestral music, increased significantly. Between 1972 and 1986, this rise in inharmonicity was accompanied by an increase in noise, but since 1986, both inharmonicity and noise have moderately decreased. In recent years (up to 2020), popular music has remained much more inharmonic than popular music from the 1960s or orchestral music involving acoustic resonance instruments. However, it has become less noisy, with noise levels comparable to those of orchestral music. We relate these trends to the evolution of music production techniques. In particular, the use of multi-tracking may explain the higher inharmonicity in popular music compared to orchestral music. We illustrate these trends with analyses of key artists and tracks.
@@ 1 Introduction
Most musical instruments that perform Western classical music are based on the acoustic resonance principle. Such instruments produce harmonic or quasi-harmonic complex tones. In contemporary popular music, the means of production are more diversified. Drums play a key role, and they do not typically produce quasi-harmonic complex tones. Instruments may be based on electronic or digital sound production techniques. The instruments’ outputs may be heavily processed with effects. Musicians may not even use instruments and produce the entirety of the music on a computer. As a result, it may not be possible to construct an efficient and appropriate description of the signal corresponding to a contemporary popular music track in terms of harmonic or quasi-harmonic complex tones.
In the spectral domain, a quasi-harmonic complex tone may be described as a series of partials whose frequencies are close to being integer multiples of one particular frequency (the fundamental). As observed by Lavengood, (2017, pp. 22, 24), there are at least two ways audio content may diverge from being quasi-harmonic: (1) the energy may not be centered around partials, in which case the signal may be described as ‘noisy’; or (2) the frequencies of the partials may not be harmonically related, in which case the tone may be described as ‘inharmonic’.
@@
One key conclusion of this study is that contemporary popular music is more inharmonic than the music from the three other datasets considered. Our diachronic analysis of the evolution of noisiness and inharmonicity in CPM leads to the conclusion that, between 1961 and 1986, recorded music evolved in a general direction that goes from acoustic tonal instruments playing content from a musical score (or, at least, that can be transcribed to a score), to heavily processed, noisy, and inharmonic content made in the recording studio and including drums. In more recent decades, music has become slightly more harmonic.
The remainder of this paper is structured as follows. Section 2 introduces the motivation for examining noise and inharmonicity in contemporary popular music, and elaborates on the relationship between the means of production and the resulting music. Section 3 forms the core of the article. It describes the datasets (Section 3.1) and the signal features used in the study (Section 3.2), before presenting a corpus-based study of inharmonicity and noisiness in CPM (Section 3.3). This section is complemented by references to other diachronic studies and remarks on the mastering process. Section 4 provides detailed analyses of specific music tracks from the perspective proposed in the core sections. The paper concludes by summarising the findings from all sections. The study is complemented by two appendices. Appendix A details the signal characteristics that influence the inharmonic character of music and discusses the link between inharmonicity and acoustic beating (or acoustical ‘grit’). It then substantiates the definitions of inharmonicity and ‘noisiness’ used in the core sections. Appendix B explains the influence of the psychoacoustic weighting used in the core sections.
@@ 2.1 On pitch and (in)harmonicity
A defining property of a harmonic tone is that it has a fundamental frequency or $f_{0}$. A classic music information retrieval task consists of estimating the fundamental frequency of a sound, a task often considered to be synonymous with ‘pitch tracking’ or ‘pitch estimation’ (Kim et al.,, 2018; Riou et al.,, 2023). The models performing such tasks are trained on harmonic or quasi-harmonic audio. For instance, Riou et al., (2023) uses MIR-1K, a dataset of quasi-harmonic samples (Hsu and Jang,, 2009) and MDB-stem-synth, a dataset of ‘exact multiples of the $f_{0}$(Salamon et al.,, 2017).
@@
Another classic music information retrieval task is audio-to-MIDI alignment (Raffel and Ellis,, 2016), in which MIDI files are aligned with the corresponding audio content. MIDI files have been referred to as symbolic ‘versions’ of pieces of music (Ewert et al.,, 2012; Raffel and Ellis,, 2016), or as ‘transcription[s]’ (Raffel,, 2016; Turetsky and Ellis,, 2003; Benetos et al.,, 2018) of the pitched content of the music. Such content has been characterised as ‘harmonic’ (Ono et al.,, 2008) and likely to be played with ‘harmonic instruments’ (Ewert et al.,, 2014).
@@
In the pitch tracking and audio-to-midi alignment tasks, pitched content is generally assumed to be harmonic or quasi-harmonic. However, closer examination of the actual practice of successful professional music producers suggests that this assumption does not hold in general—at least for contemporary popular music. For example, Figure 1 shows the power spectrum of one bass ‘note’ in the 2023 track ‘¡Fire!’ by Primaal (Primaal,, 2023). We write ‘note’ in scare quotes, as its frequency content is not at all harmonic, even though it provides a strong impression of pitch. This example is not unique—there are many such examples of sounds in Primaal’s productions that evoke a sense of pitch, while being highly inharmonic. Primaal is a brand of the Hyper Music production company (Hyper Music,, 2024) mentioned by Deruty et al., (2022). In 2022 and 2023, Primaal authored music for commercials commissioned by brands such as L’Oréal, Adidas, Vichy, Honda, GoPro and Chanel. The prominence of these brands and the great size and diversity of the intended audience for this music suggests that Primaal’s productions might reasonably be considered to represent important current trends in the music industry.
@@
Figure 1: Power spectrum for the first bass ‘note’ in Primaal’s ‘¡Fire!’. The vertical red and blue lines represent the fundamental and expected positions of the harmonics for the two highest spectral peaks.
@@
There therefore seems to be a discrepancy between such examples of the use of inharmonicity in real-world music production and the standard view within music information retrieval that pitched content corresponds to harmonic tones. This discrepancy provides motivation for exploring the extent to which pitch-evoking sounds in CPM are harmonic.
2.2 On the influence of technology on music
The example in Figure 1 illustrates how some of the resulting sounds in CPM can only be achieved using specific technologies. The example features audio produced by a composite setup of instruments: a Spectrasonics Omnisphere bass superimposed on a Eurorack setup using FM synthesis, resonators, and ring modulators. Producing this with acoustically resonant instruments would have been challenging, if not practically impossible.
This section explores the broader relationship between technology and music. In Section 2.2.1, we discuss the link between performance resources and musical output. Section 2.2.2 briefly covers modern music production technology. In Section 2.2.3, we explain why we have focused on inharmonicity and noise, particularly as CPM may be moving away from using musical notes as primary structuring elements.
2.2.1 Resources of the performance and musical idioms
Music involves what Parry, (1911) calls the resources of the performance, such as the human voice and musical instruments. In Western classical music, many instruments are based on acoustic resonators, like vibrating strings or air columns, which produce waveforms with strong periodicity. These waveforms are often modelled as the sum of components based on a fundamental frequency and its overtones, which are (close to) integer multiples of that fundamental (Young,, 1952).
According to Huron and Berec, (2009, p. 103), idiomaticism is, ‘of all the ways a given musical goal or effect may be achieved, the method employed by the composer/musician [that] is one of the least difficult’. For instruments with acoustic resonators, producing harmonic or quasi-harmonic sounds is easy, making these sounds idiomatic to such instruments.
Such sounds have formed the basis of much musical discourse. In acoustic resonators, pitch impressions from harmonic or quasi-harmonic sounds are perceived as musical notes, which are used to construct phrases and chords. According to Pascall,’s (2001) terminology, the idiom lies in the musical phrases and chords enabled by the instrument. The properties of resonator-produced sounds can shape musical discourse. For example, harmonic positioning influences musical temperament (Barbour,, 2004), and Western counterpoint rules, such as avoiding parallel unisons, octaves, and fifths, relate to the waveform properties of resonators (Huron,, 2001).
Technological progress expands musical idioms by providing more resources. In the 20th century, signal analysis allowed musicians like Gérard Grisey to create spectral music, where scores derive from harmonic positions of acoustic instruments, with deviations creating a gradation from harmonic to inharmonic textures (Rose,, 1996). Here, the idiom in Huron and Berec, (2009)’s sense is orchestral texture, while in Pascall, (2001)’s sense, it is gradation.
The 20th and 21st centuries saw major innovations in music technology (Webster,, 2002), such as the Ondes Martenot (1928) (Orton,, 2001), the Moog synthesiser (1964) (Porcaro,, 2013), and the FM-based Yamaha DX7 (1983) (Mattis,, 2001), pivotal in ’80s popular music (Lavengood,, 2017, p. 36). Sound recording technology, evolving since 1877 (Mumma et al.,, 2001), also impacted musical idioms. For instance, continuously shifting frequencies are idiomatic to the Ondes Martenot (Huron and Berec,, 2009), while distortion, linked to analog circuitry, defines genres like heavy metal.
2.2.2 The ‘utopian vision of a boundless space for musical exploration’
Technological innovation affects more than just the advent of specific idioms. Music-related technology takes various forms and influences different stages of music production and engagement. One such area is the compositional process. For example, before the 16th century, paper was expensive, and erasable pencils did not exist (Charlton and Whitney,, 2001), limiting composers’ ability to draft and sketch musical ideas (Bent,, 2001). The introduction of cheaper paper and pencils transformed composition.
More recently, technological progress has allowed the recording studio to become a creative tool in its own right (Moylan,, 2014). Spicer, (2004) links ‘accumulative’ and ‘cumulative’ forms in pop/rock music to advances in recording technology. The 1990s saw the rise of home studios, making professional practices accessible. Almost anyone with a computer could assume the role of a musician–engineer hybrid (Pras et al.,, 2013; Bell,, 2014), with a focus on micro-manipulating digital audio (Théberge,, 2001).
We argued that using musical notes is idiomatic to acoustic resonators, which impose the constraint (McPherson and Tahıroğlu,, 2020) that music involves ‘notes’—symbols representing harmonic or quasi-harmonic tones based on a fundamental frequency that stays constant during the note’s duration. In the studio, this constraint vanishes. A key question is what idiom musicians use without this constraint: ‘even if we were to achieve the utopian vision of a boundless space for musical exploration, we would still be left with the question of what possibilities musicians would choose to explore within it’ (McPherson and Tahıroğlu,, 2020).
2.2.3 To what extent do idioms of popular music involve musical notes?
Musique concrète exemplifies music where musicians have abandoned the musical note as a building block, working in the recording studio (Schaeffer,, 2020). This genre, derived from Schaeffer,’s (1966) experiments, often lacks clear pitch sensations, yet it is not devoid of pitch. As Yost, (2009) states, ‘[m]usic without pitch would be drumbeats, speech without pitch processing would be whispers’. This does not entirely apply to musique concrète, as seen in works like Pierre Henry’s Temps de Pointe or Bernard Parmeggiani’s Ondes. Nonetheless, transcribing it into standard notation is difficult, and instead, transcriptions may involve free shapes representing dynamics or timbre (Favreau et al.,, 2010).
From a commercial perspective, musique concrète is niche. The music charts indicate that the music that is most listened to is recent popular music. Popular music is distinguished from folk and art music by its use of recorded sound as the main mode of transmission (Tagg,, 1982; Middleton,, 1990; Mazzanti,, 2019). It is characterised by commercial interests, entertainment, and ties to mass media (Frith,, 2004).
Deruty et al., (2022) define Contemporary Popular Music (CPM) as current popular music, known for its technological innovation and cross-genre influences (Mazzanti,, 2019; Frith,, 2004). CPM genres include post-rock, rap/hip-hop, electronica, and non-Western styles like K-pop and reggaeton. Bertin-Mahieux et al., (2011) use the term, contemporary popular music, to refer to tracks from 1922 onwards. We use it to refer to popular music since the 1970s, with Sgt. Pepper’s Lonely Hearts Club Band (McCartney and Miles,, 1997) as a milestone marking the beginning of the period. Like musique concrète, recent popular music is studio-produced. In Parry,’s (1911) terms, both genres share the same ‘resources of the performance’ (the studio), with similar constraints and affordances.
Has popular music abandoned musical notes? Huang et al., (2020) identify building blocks in popular music. The ‘melody’ and ‘harmony’ blocks suggest that popular music can be transcribed to musical notes, either automatically (Bertin-Mahieux et al.,, 2011) or manually (Tagg,, 1982). On the other hand, percussion, important enough for research in automatic transcription (Wu et al.,, 2018) and musicology (Mowitt,, 2002), is characterised by clear stochastic noise components in instruments like snare drums (Fletcher and Rossing,, 2012, pp. 602-606), making their transcription to musical notes difficult.
Percussion is not the only aspect of CPM that complicates transcription to musical notes. For example, sampling introduces real-life noisy sounds into music (Forman and Neal,, 2004, p. 408), and the noise surrounding each partial in guitar distortion (Berger and Fales,, 2005, p. 184) can blur frequencies, making pitch identification difficult. The fact that not all popular music can be transcribed meaningfully into notes (i.e., staff notation) raises the question: can we quantify how much popular music deviates from being a combination of ‘notes’? In other words, to what extent does it deviate from being produced (or producible) by acoustic resonators or technology that emulates them? This article aims to begin answering this question. We hope this article provides the beginning of an answer to this question.
3 Datasets, features, analysis
@@ 3.1 Datasets
    The popular music dataset, or ‘BEA dataset’, contains 30,435 tracks released between 1961 and 2022, with at least 460 tracks per year. It extends the dataset used by Deruty and Pachet, (2015), and is based on the ‘Best Ever Albums’ website (Best Ever Albums,, 2024), which aggregates reviews to rank top-rated albums each year. As these ratings are dynamic, the dataset reflects a snapshot at the time of writing. Both original and remastered versions are included, and the influence of remastering is discussed in Section 3.5.
@@
    The orchestra dataset contains 10800 orchestral and opera tracks from various sources, including Deutsche Grammophon’s ‘Classical Gold’, ‘The History of Classical Music’, Decca’s ‘55 Great Vocal Recitals’ and ‘Ultimate Boxset’ series. Tracks in this dataset range from the Baroque era to the early 20th century. Different interpretations of the same original score may be found in the dataset.
@@
    The musique concrète dataset consists of 1000 tracks from composers related to Pierre Schaeffer’s school of thought by either having directly collaborated with him or having produced music at INA/GRM in Paris. It includes music by Pierre Henry, Bernard Parmegiani, Denis Dufour, and François Bayle amongst others.
@@ 3.2 Features
One key aspect of the study is the quantitative evaluation of inharmonicity in music. Inharmonicity has been defined as ‘the departure in frequency from the harmonic modes of vibration’ (Young,, 1952) and ‘the deviation of a set of frequencies from an exact harmonic series’ (Campbell,, 2001). A further distinction can be observed between ‘inharmonic sounds which have little if any relevance for music (e.g., white or pink noise)’ (Schneider and Frieler,, 2009) and ‘coherent’ inharmonic signals, which ‘sound as stable and smooth as harmonic signals’ (De Boer,, 1956). In other words, inharmonicity can be the result of either inharmonic partial relations or noise (Appendix A.1.6 elaborates on the relationship between noise and inharmonicity). As a result, the features we use must address both partial relations and noise.
@@ 3.2.1 HR-inharmonicity
We use the Matlab R2022b implementation of the MPEG-7 feature HarmonicRatio as part of the broader inharmonicity analysis. [^2] https://fr.mathworks.com/help/audio/ref/harmonicratio.html HarmonicRatio has been described as measuring ‘the proportion of harmonic components within the power spectrum’ (ISO,, 2001, p. 36) and ‘the ratio of harmonic power to total power’ (Moreau et al.,, 2006, p. 33). To our knowledge, HarmonicRatio is the only feature in the literature that measures the proportion of harmonic components in the general case. Features such as Inharmonicity in Peeters, (2004, p.17) measure ‘the […] divergence from a purely harmonic signal’ and can only be measured in relation to a single complex tone.
HarmonicRatio derives from the normalised auto-correlation of the signal. The normalisation is performed so that the auto-correlation at zero lag equals one. The feature output is the maximum value of the normalised auto-correlation after the first zero-crossing. HarmonicRatio assumes values between 0 and 1, with harmonic complex tones resulting in 1. As we want to measure inharmonicity, we define a feature called HR-inharmonicity, defined as $1-HarmonicRatio$. Appendix A.1 elaborates on the relationship between different aspects of the signal and HR-inharmonicity.
Patterson et al., (1996) and Yost, (1996, 1997) have argued that the ‘first peak of the auto-correlation function’ can be used to evaluate the ‘saliency or the strength of pitch of complex sounds’—in other words, the perceptual impression of pitch strength. The matter was further investigated by Shofner and Selas, (2002). The illustrations in Yost, (1996, p. 3330) and Shofner and Selas, (2002, p. 439) show that the ‘first peak of the auto-correlation function’ is the maximum value of the normalised auto-correlation after the first zero-crossing—that is, the same as what the MPEG-7 standard designates as HarmonicRatio (ISO,, 2001; Moreau et al.,, 2006). One of our main concerns in this study is quantifying the degree to which a signal deviates from being harmonic. However, as HarmonicRatio can be used to gauge the strength of pitch, the properties of the signal that we find to result in lower (or higher) HarmonicRatio values may also lead to a corresponding decrease (or increase) in the perceived strength of pitch.